A motor-cyclist travels at a constant speed of 72 kmh-1 along a straight tevel road towards an

observer standing at the center of the bridge over the road. When the motor- cyclist is distant, the

observer hears a sound of frequency 65,0 Hz from the engine of the motor-cycle. What would be

the frequency of the sound heard by the observer after the motor cycle has passed under the bridge

and is along way from the observer?



1
Expert's answer
2021-03-17T16:58:22-0400
"f_o=\frac{c}{c-v_s}f_s\\65=\frac{340}{340-\frac{72}{3.6}}f_s\\f_s=61.2\ Hz"